OpenStudy (inkyvoyd):

∫ sec(x) dx=? Please just guide me along, because I want to get that sense of accomplishment of figuring something out. ∫ sec(x) dx -> ∫ (sec(x)tan(x))/(tan(x) dx u=sec x du/dx=sec x tan x dx = du/(sec x tan x) --> ∫u/u' du Then, integration by parts f(x)=u, g'(x)=1/u' u ln u' - ∫ 1*ln u' du =u ln u' -u' ln u' - u' =(u-u')ln u' -u' resubstitute (sec x-sec x tan x)ln(sec x tan x)- sec x tan x I know I forgot the abs signs some where, and I'm wondering if I even did this right.

OpenStudy (inkyvoyd):

Also, thanks a lot Wolfram Alpha, for not explaining this at all. -_-

OpenStudy (amistre64):

multiply by a useful form of "1" sec+tan ------ sec+tan

OpenStudy (amistre64):

one issue i see that might be throwing you for a loop is that sec tan is not defined in us

OpenStudy (anonymous):

Int Secx => you have to multiply Secx by \[(\sec x+\tan x)/(\sec x+\tan x)\] in order to have the derivative of the bottom on top. You should get: \[\int\limits_{?}^{?}(\sec ^{2}x+\sec x \tan x)/(\sec x+\tan x)\] Now you can just do u-sub u=secx + tanx du = (sec x tan x +Sec^2)dx You will end up with:\[\int\limits_{?}^{?}\sec x dx=\ln |\sec x+\tan x|+C\]

OpenStudy (amistre64):

you havent learned to type in latex yet i see

OpenStudy (anonymous):

Ha no, too complicated

OpenStudy (inkyvoyd):

It was sec x+tan x?! FUUUUUUUUUUUUUUUUU

OpenStudy (amistre64):

\int for \(\int\) \frac{n}{d} for \(\cfrac{n}{d}\)

OpenStudy (inkyvoyd):

Wait, is my integration correct though? I know it was a ridiculously convoluted way, but...

OpenStudy (amistre64):

i dont think so; you introduced a flaw in it at the beginning i think that persisted towards the end

OpenStudy (amistre64):

dx = du/sec(x) tan(x) would be fine ... but youd have to express tan(x) in terms of u as well